40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Sep 19 01:49:52 2020, Koji, Update, General, M4.5 EQ in LA 
    Reply  Sat Sep 19 11:27:04 2020, rana, Update, General, M4.5 EQ in LA Sep18-EQ.pdf
       Reply  Sat Sep 19 18:08:34 2020, rana, Update, General, M4.5 EQ in LA 
    Reply  Sat Sep 19 19:14:59 2020, Koji, Update, General, ITMX released / ETMY UR magnet knocked off IMG_6218.jpeg
       Reply  Sun Sep 20 11:41:54 2020, gautam, Update, General, IMC re-locked 
       Reply  Mon Sep 21 12:40:58 2020, gautam, Update, General, Workable IFO recovery IFOrecovery.pngIMG_5345.JPG
Message ID: 15580     Entry time: Sat Sep 19 01:49:52 2020     Reply to this: 15581   15585
Author: Koji 
Type: Update 
Category: General 
Subject: M4.5 EQ in LA 

M4.5 EQ in LA 2020-09-19 06:38:46 (UTC) / -1d 23:38:46 (PDT) https://earthquake.usgs.gov/earthquakes/eventpage/ci38695658/executive

I only checked the watchdogs. All watchdogs were tripped. ITMX and ETMY seemed stuck (or have the OSEM magnet issue). They were left tripped. The watchdogs for the other SUSs were reloaded.

ELOG V3.1.3-